find the volume of the solid generated by revolving region bounded by the curves y = x, y = 2 - x, y = 0, about y = -1


1
Expert's answer
2022-01-31T13:29:55-0500
"x=2 - x=>x=1"

"x=0, 2-x=0=>x=2"

"V=\pi\displaystyle\int_{0}^{1}((x+1)^2-(0+1)^2)dx"

"+\pi\displaystyle\int_{1}^{2}((2-x+1)^2-(0+1)^2)dx"

"=\pi[\dfrac{x^3}{3}+x^2]\begin{matrix}
 1 \\
 0
\end{matrix}+\pi[8x-3x^2+\dfrac{x^3}{3}]\begin{matrix}
 2 \\
 1
\end{matrix}"

"=\pi(\dfrac{1}{3}+1+16-12+\dfrac{8}{3}-8+3-\dfrac{1}{3})"

"=\dfrac{8\pi}{3}({units}^3)"